OpenStudy (anonymous):

If f(x) = 0.5(2 x), what is the value of f(2)? (Points : 5) 0 0.5 2 4.5

Mehek (mehek14):

put 2 for x

OpenStudy (anonymous):

oh

Mehek (mehek14):

\(f(x) = 0.5(2*2)\)

Mehek (mehek14):

\(2*2 = 4\)

Mehek (mehek14):

\(0.5*4=2\)

OpenStudy (anonymous):

oh thank you i understand most of it

Mehek (mehek14):

All you have to do is put 2 for x and then solve the equation

OpenStudy (anonymous):

thank you very much

Mehek (mehek14):

Yw